OpenStudy (anonymous):

someone help me with this. im so confused! Tara's car travels about 22 miles per every one gallon of gas. Her tank holds anywhere from 18-22 gallons of gas. what is a reasonable domain and range for the function that relates miles traveled to the amount of time she has to fill up her tank?

OpenStudy (triciaal):

travels 22 miles per gallon lowest amount = 18 gallons lowest distance = 22 miles per gallon for 18 gallons = (22* 18 ) miles= 396 miles same rate 22 mpg highest/maximum gallons = 22 distance =22* 22 miles- 484 miles (22 - 18) *22 = 88 miles we know how far Tara can travel on a full tank 396 to 484miles I don't see anything relating the time. Did you copy the question correctly?

OpenStudy (anonymous):

amount of times she has to fill up her tank?*

OpenStudy (triciaal):

what distance is she planning to travel ? at what speed/velocity is she travelling?